Circular No. 8129

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003ed IN NGC 5303A S. Nakano, Sumoto, Japan, reports the discovery by Koichi Itagaki

(Teppo-cho, Yamagata-ken) of an apparent supernova (mag 15.2) on unfiltered CCD frames taken around May 3.496 and 4.479 UT. An image taken by Itagaki with a 0.60-m reflector on May 3.526 yields the following precise position for SN 2003ed: α = 13^{h}47^{m}45^{s}.36, δ = +38^{o}18′20′′.3 (equinox 2000.0), which is 8′′ northeast of the nucleus of NGC 5303. Nothing appeared at this location on images taken by Itagaki on Apr. 9 and 27 (limiting mag 19). An unfiltered CCD image taken by R. Kushida, Yatsugatake South Base Observatory, on May 3.604 (measured by Y. Kushida) yields position end figures 45^{s}.38, 20′′.2 (mag 15.3), which is 3′′.3 east and 4′′.1 north of the center of NGC 5303A (with NGC 5303B situated southwest of NGC 5303A). K. Kadota, Ageo, Japan, reports position end figures 45^{s}.39, 20′′.8 (and mag 15.2) for SN 2003ed from an unfiltered CCD image taken on May 3.617.

T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2003ed (cf. IAUC 8129), obtained by J. Huchra on May 5.32 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova soon after explosion. The spectrum consists of a blue continuum with incipient P-Cyg lines of hydrogen and helium.

SUPERNOVAE 2003dt AND 2003dw Matheson et al. add that a spectrum of SN 2003dt (cf. IAUC 8120),

obtained as above on May 5.45 UT, shows it to be a type-Ia supernova several weeks past maximum. A spectrum of SN 2003dw (cf. IAUC 8124), obtained on May 5.34, shows it to be a type-Ia supernova near maximum. Adopting the NASA/IPAC Extragalactic Database recession velocity of 9003 km/s for the host galaxy, the supernova expansion velocity is 12500 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova (Riess et al. 1997, A.J. 114, 722) is 1 ± 2 days after maximum light.

SUPERNOVA 2003ds IN MCG +08-19-17 Corrigendum. On IAUC 8120, line 8, for SN 2003ds in NGC 3191 read

SN 2003ds in MCG +08-19-17

2003 May 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green